A full grown giraffe is about 6 times heavier than a full grown lion. A full grown cheetah is about one quarter as heavy as a full grown lion. together, they are about 450kg. what is the approximate mass of each animal.
Answer by stanbon(57410)

A full grown giraffe is about 6 times heavier than a full grown lion. A full grown cheetah is about one quarter as heavy as a full grown lion. together, they are about 450kg. what is the approximate mass of each animal.
---------------
Let the lion's weight be "x" kg
Then the giraffe's weight is "6x" kg
And the cheetah's weight is (1/4)x kg
-------------------
EQUATION:
x+6x+(1/4)x = 450 kg
(7 1/4)x= 450
7.25x = 450
x= 450/7.25 = 62.07 kg (approx weight of lion)
6x=372.41 kg (approx weight of giraffe)
(1/4)x = 15.51 kg (appox weight of cheetah)
====================
Cheers,
Stan H.